OpenStudy (anonymous):

How do you factor 4x2 – 4x – 48 ?

Parth (parthkohli):

Hey there! Factoring out the 4 is a good start. \( \color{Black}{\Rightarrow 4(x^2 - x - 12) }\)

Parth (parthkohli):

If you notice, -4x + 3x = -x -4x * 3x = -12x Split the middle term. \( \color{Black}{\Rightarrow 4(x^2 - 4x + 3x - 12x) }\)

Parth (parthkohli):

And then you just group. \( \color{Black}{\Rightarrow 4(x(x - 4) + 3(x - 4)) }\) \( \color{Black}{\Rightarrow 4(x +3)(x - 4)}\) Did you get it? 0mgl0la

OpenStudy (anonymous):

It is quite simple 4(x^2-x-12) =4(x^2-4x+3x-12) =4x((x-4)+3(x-4)) =(4x-3)(x-4)

OpenStudy (anonymous):

@ParthKohli Can you please look at the questions I have posted need help exam on 9th

Parth (parthkohli):

I answered that to you in chat, @ipm1988

OpenStudy (anonymous):

4x^2-4x-48 (2x+6)(2x-8).... :-)

OpenStudy (anonymous):

So, what are the factors..?

Parth (parthkohli):

Hey! Scroll up :)

OpenStudy (anonymous):

(4x-3)(x-4) ?

Parth (parthkohli):

And, @ipm1988 I believe that it is (4x + 3)(x - 4) ;)

Parth (parthkohli):

Actually, it is \(4(x + 3)( x- 4)\)

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

@ipm1988 ,,can u get -4x in that factor??i think we can't.. just confusing

OpenStudy (anonymous):

@ParthKohli you are right i made a mistake sorry i was not able to see your message in chat what did you wrote

Parth (parthkohli):

I am just 12, so I can't do Calculus :(

OpenStudy (anonymous):

no that's ok,,i'm just confuse by your answer,, :)

OpenStudy (anonymous):

my answer is wrong